OpenStudy (anonymous):

A piece of wire 5 inches long is to be cut into two pieces. One peice is x inches long and is to be bent into the shape of a square. The other piece is to be bent into the shape of a circle. Find an expression for the total adrea made up by the square and the circle as a function of x.

OpenStudy (aum):

The wire is 5 inches long. It is cut into two pieces. The first piece is "x" inches. Therefore, the second piece is (5-x) inches. If you make a square out of a wire that is "x" inches long, what will be the length of each side?

OpenStudy (aum):

A square has sides of equal lengths. Therefore, an "x" inch wire bent into the shape of a square will have each side x/4 inches long. What is the area of a square when each side is x/4 inches long?

OpenStudy (jack1):

you're doing awesome so far 5 = x + y x = perimeter of square y = circumference of circle \(\Large \text {Area of square} = (\frac {\text perimeter}{ 4})^2 \) \( \Large= (\frac x4)^2\) \( \Large= \frac x{16}\)

OpenStudy (jack1):

\(\Large \text {Area of Circle} = (\pi \times r)^2 \) \(\Large Circumference = 2 \times \pi \times r \) \(\Large y = 2 \times \pi \times r \) \(\Large (5-x) = 2 \times \pi \times r \) \(\huge \frac {(5-x)}{2 \pi} = r \) sub this into the above equation for area of a circle then total area = sum of both of these shapes area